OpenStudy (anonymous):

The sides of a triangle have measures of x, x, and y, and the perimeter of the triangle is 46. The measure y is 4 centimeters longer than the measure x. Write the solution for this as an ordered pair.

OpenStudy (anonymous):

asdasd

OpenStudy (anonymous):

erm. so 2x+y =46 2x+x+4=46 (sub in the y as x+4 because it tells you that) 3x=46-4 x=42/3

OpenStudy (anonymous):

and hi (:

OpenStudy (anonymous):

Haha hey! (: lol and thanks =3

OpenStudy (anonymous):

>: dont remember me? O: QQ

OpenStudy (anonymous):

Of course i do!!!! froggy!!!! :p

OpenStudy (anonymous):

:3 anyway just bug me for anything else, you can post it here. idc

OpenStudy (anonymous):

Haha and you thought i didnt remember u.... tsk tsk.... :D haha and okay... hold on (:

OpenStudy (anonymous):

Waiit... it says ordered pair...:P

OpenStudy (anonymous):

yup so solve for the x then do (x,x+4) is my guess of what they want .-. So then it's (x,y) which is what an ordered pair is ?

OpenStudy (anonymous):

Yep lol Okay so then hold on a secy.... ^^

OpenStudy (anonymous):

haha so then its just (14,18) Correcto? :D

OpenStudy (anonymous):

sounds right

OpenStudy (anonymous):

double check your answer to see if it's right. I don't get how you solved this. But if u are not sure. just double check. it won't hurt to do it.

OpenStudy (anonymous):

Haha okay (: